Eve was twice as old as her sister Teresa in 2006. Eve was born in 1992. How old will Teresa be in 2009.
Can you please help me, I am lost when comes to Math.
Please provide a step by step solution
Appreciate it!
Thanks


Found 2 solutions by edjones, checkley75:
Answer by edjones(8007)

Do NOT panic!!!
You CAN do this.
2006
1992
------subtract
0014
.
So, when eve was 14 she was twice as old as her sister.
How old was her sister? What is 14 divided by 2? Yes, 7.
So, if she was 7 in 2006, how old will she be in 2009? Yes, 10!!!
.
If you don't give up and try these problems you will start to get some.
After a while you will get more. Work hard and you could become a whizz!!!
.
Ed

Answer by checkley75(3666)

E=2T IN 2006
E=2006-1992=14 YEARS OLD.
14=2T
T=14/2=7 YEARS OLD IN 2006
T WILL BE [7+(2009-2006)=7+3=10] YEARS OLD IN 2009.